# One-loop QCD corrections to SSA in unweighted Drell-Yan processes

### v1

@@ V.1.2 Non-pole contribution
For non-pole contribution, $x\neq 0$, and only the real part of $B(s_{i})$ contributes to $g_{i}$. The combination $B(s_{2})-B_{c}(s_{2})$ is purely imaginary and does not contribute. So, $g_{0,1,2}$ are zero. Moreover, the difference between $D_{3}$ and $D_{4}$ is proportional to $B(s_{2})-B_{c}(s_{2})$, thus the non-pole part of $g_{3}$ and $g_{4}$ is the same, i.e.,
@@
$\frac{1}{N_{c}C_{F}}p_{a}^{+}C^{NP}_{3}=$ $-\frac{4x_{2}C_{A}}{x_{1}s_{2}}\text{Re}\Big{(}B(s_{0})-B(s_{2})\Big{)}-\frac{4x_{2}}{xs_{2}}\text{Re}\Big{(}C_{A}B(s_{0})-(C_{A}-2C_{F})B(s_{2})\Big{)}$
$-\frac{8(C_{A}-2C_{F})}{\epsilon}\frac{x_{2}^{2}}{s_{2}x^{2}}\text{Re}\Big{(}B(s_{1})-B(s_{2})\Big{)}.$ (88)
Because $s_{0},s_{1}$ can be negative or positive, $C_{3}^{NP}$ is non-zero. Note that
$\frac{1}{N_{c}C_{F}}p_{a}^{+}\Big{[}\frac{1}{x_{2}}C_{3}+\frac{1}{x_{b}}D_{3}\Big{]}=$ $\frac{8C_{F}}{s_{2}(x+i\epsilon)}\frac{4-\epsilon+\epsilon^{2}}{\epsilon}(B(s_{2})-B_{c}(s_{2})),$ (89)
which does not contain non-pole contribution. So, the non-pole parts of $C_{3}$ and $D_{3}$ satisfy
@@
Both $t_{a}^{\mu\nu}$ and $t_{b}^{\mu\nu}$ appear in this part, while only $t_{a}^{\mu\nu}$ appears in SGP part and in tree level result. Moreover, the coefficients here are divergent. Different from twist-2 cases, the new tensor structure with divergent coefficients does not imply a breaking of factorization, as we will explain later.
@@ V.2 Corrections to derivative part of $W^{\mu\nu}$
$H_{6}^{\mu\nu}=H_{4}^{\mu\nu}=-i\frac{\partial\delta^{n-2}(q_{\perp})}{\partial q_{\perp}^{\rho}}\tilde{s}_{\perp}^{\rho}Tr[H^{\mu\nu}\otimes T^{a}\gamma^{-}].$ (93)
@@
which is just the correction to quark form factor, as pointed out in Chen et al. (2017). Note that $s_{2}=Q^{2}$ for virtual correction.
@@ V.3 Total virtual corrections
Now, the complete virtual correction to hadronic tensor is the sum of eqs.(86,91,94), that is,
@@
If there is no non-pole contribution, the virtual correction has the same structure as tree level hadronic tensor. The pole contribution, the last second line of above result, satisfies QED gauge invariance, as shown for tree level hadronic tensor. The correction to derivative part is the same as quark form factor, in agreement with the conclusion of Chen et al. (2017). In Chen et al. (2017), such correction is inferred from Ward identity for longitudinal gluon $G^{+}$. Here we recover the result by direct calculation. The non-pole contribution, the last line, also satisfies such invariance: with $q_{\perp}=0$,
@@
From $W^{\mu\nu}$, $I\langle L\rangle$ and $I\langle P_{7}\rangle$ can be obtained easily,
@@
$I\langle L\rangle\Big{|}^{NP}_{v}=$ $2\times l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int\frac{d\xi}{\xi}\frac{dx_{b}}{x_{b}}\frac{dx_{2}}{x_{2}}\bar{q}(x_{b})T_{F}(x^{*}_{2}-x_{2},x^{*}_{2})\delta(1-\hat{x}_{b})$
$\times\Big{(}128\hat{x}_{2}\Big{[}(\frac{1}{\epsilon}+1)\theta(x_{2}-x^{*}_{2})-\theta(x^{*}_{2}-x_{2})\frac{1}{2}\ln\frac{\hat{x}_{2}}{\hat{x}_{2}-1}\Big{]}$
$-\theta(-x_{2})\frac{64N_{c}^{2}}{1-\hat{x}_{2}}+64(1-N_{c}^{2}\theta(x_{2}))\Big{)}.$ (100)
@@
$\hat{x}_{2}\equiv\frac{x_{2}^{*}}{x_{2}},\ \hat{x}_{b}\equiv\frac{x_{b}^{*}}{x_{b}},\ x_{2}^{*}=\xi,\ x_{b}^{*}=\tau_{\xi},\ \xi=\frac{q^{+}}{p_{a}^{+}}.$ (102)
@@ VI Real corrections
For real corrections, $q_{\perp}\neq 0$ is assumed. $W^{\mu\nu}$ thus depends on $p_{a}^{\mu},p_{b}^{\mu},q_{\perp}^{\mu}$ and $\tilde{s}_{\perp}^{\mu}$. Because of QED gauge invariance $q_{\mu}W^{\mu\nu}=0$, there are six independent tensor structures for transversely polarized caseArnold et al. (2009). For our convenience, the tensors can be chosen as
@@
The dot product represents the contraction of Lorentz indices, e.g., $P_{j}\cdot W=P_{j,\mu\nu}W^{\mu\nu}$. Eq.(105) can be solved directly. But the coefficients before $P_{i}\cdot W$ may depend on spin vector $\tilde{s}_{\perp}$ in a very complicated way, since $P_{i}\cdot P_{j}$ depends on $\tilde{s}_{\perp}$. Noticing that $W_{i}$ does not depend on the direction of $q_{\perp}$, we can integrate out the angles of $q_{\perp}$ on both sides of eq.(105) and then to solve the obtained equations. According to this treatment, we successfully get
$W_{i}=\frac{2-\epsilon}{\Omega_{n-2}\tilde{s}_{\perp}^{2}}C_{ij}\xi_{j},\ \xi_{j}\equiv\int d\Omega_{n-2}P_{j}\cdot W,$ (106)
@@
Notice that $C_{ij}$ is symmetric, $C_{ij}=C_{ji}$. $q_{t}=|\vec{q}_{\perp}|$.
@@
$=$ $\sum_{i}\int d^{n}q\delta(q^{2}-Q^{2})L\cdot P_{i}\frac{2-\epsilon}{\Omega_{n-2}\tilde{s}_{\perp}^{2}}C_{ij}\xi_{j}.$ (108)
@@
$\Big{(}\int d\Omega_{n-2}L\cdot P_{i}\Big{)}\frac{2-\epsilon}{\Omega_{n-2}\tilde{s}_{\perp}^{2}}C_{ij}\xi_{j}=l_{\perp,cs}\cdot\tilde{s}_{\perp}\sum_{i}a_{i}\xi_{i}\frac{1}{\tilde{s}_{\perp}^{2}}.$ (111)
@@
This is our main formula for real corrections. There is no $\tilde{s}_{\perp}$ or $l_{\perp,cs}$ in $\tilde{W}_{i}$ now. With $W^{\mu\nu}$ replaced by $\tilde{W}_{i}$, the formula eq.(40) can be applied. Further calculation is the same as that for real correction of weighted cross section studied in Chen et al. (2017).
Same as weighted cross section, there is no contribution from two-point correlation functions, i.e.,Fig.3. The amplitudes in these diagrams do not contain any absorptive part. So, $M^{(0)}_{\gamma_{\perp}\gamma_{5}}$ does not contribute and we just need to consider three-point distribution functions. Still, three types of poles contribute: SGP, SFP and hard pole(HP). The details can be found in Chen et al. (2017), here we just present the final result for each pole contribution. We note that for HP and SFP, the momentum fraction of initial gluon $x\neq 0$, and it can be shown easily that the collinear expansion based on $G^{+}$ or $G_{\perp}$ leads to the same hard coefficientsEguchi et al. (2007). We have checked this by direct calculation. In App.B, we present some details for the calculation of HP contribution based on $G_{\perp}$ expansion. The procedure based on $G^{+}$ expansion is too lengthy and not shown. For SGP contribution, we use $G^{+}$ to do collinear expansion only. Especially, for Fig.4, the contribution from $M^{(1)}_{\gamma_{\perp}\gamma_{5}}$ vanishes, because corresponding $W^{\mu\nu}$ projected by $\gamma_{5}\gamma_{\perp}$ is anti-symmetric in $\mu,\nu$. $M^{(1)}_{\gamma_{\perp}}$ does not contribute because
@@
which can be shown from PT symmetry or from the second equation of eq.(25) and eq.(31). In this work, we consider only the contribution proportional to $\bar{q}\otimes T_{F}$. Because $M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}$ and $M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}$ are related to $T_{\Delta}$ rather than $T_{F}$, we do not consider the contribution from these two matrix elements in the following. Our SGP contributions are given by $M^{(1)}_{\gamma^{+},\partial_{\perp}\psi}$ and $M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}$. As illustrated in Sec.III, SGP contributions from these two matrix elements can be expressed by $T_{F}$. In addition, we also use $L^{\mu\nu}$ instead of the projection operators $P_{i}^{\mu\nu}$ to calculate. The same $I\langle L\rangle$ is obtained. This is a check of our calculation.
@@
VI.1 SGP contribution
SGP contribution is given by Fig.4. The short bar indicates the propagator is on-shell. Fig.4(c,d) are mirror diagramsEguchi et al. (2007), which do not contribute. Our calculation confirms this.

### v2

@@ V.1.2 Non-pole contribution
For non-pole contribution, $x\neq 0$, only the real part of $B(s_{i})$ contributes to $g_{i}$. The combination $B(s_{2})-B_{c}(s_{2})$ is purely imaginary and does not contribute. So, $g_{0,1,2}$ are zero. Moreover, the difference between $D_{3}$ and $D_{4}$ is proportional to $B(s_{2})-B_{c}(s_{2})$, thus the non-pole part of $g_{3}$ and $g_{4}$ is the same, i.e.,
@@
$\frac{1}{N_{c}C_{F}}p_{a}^{+}C^{NP}_{3}=$ $-\frac{4x_{2}C_{A}}{x_{1}s_{2}}\text{Re}\Big(B(s_{0})-B(s_{2})\Big)-\frac{4x_{2}}{s_{2}(x)_{p}}\text{Re}\Big(C_{A}B(s_{0})-(C_{A}-2C_{F})B(s_{2})\Big)$
$-\frac{8(C_{A}-2C_{F})}{\epsilon}\frac{x_{2}^{2}}{s_{2}(x^{2})_{p}}\text{Re}\Big(B(s_{1})-B(s_{2})\Big).$ (93)
$1/(x)_{p}$ is for the principal value of $1/x$. $1/(x^{2})_{p}$ is well defined because $B(s_{1})-B(s_{2})=0$ when $x=0$. Because $s_{0},s_{1}$ can be negative or positive, $C_{3}^{NP}$ is non-zero. Note that
$\frac{1}{N_{c}C_{F}}p_{a}^{+}\Big[\frac{1}{x_{2}}C_{3}+\frac{1}{x_{b}}D_{3}\Big]=$ $\frac{8C_{F}}{s_{2}(x+i\epsilon)}\frac{4-\epsilon+\epsilon^{2}}{\epsilon}(B(s_{2})-B_{c}(s_{2}))$ (94)
does not contain non-pole contribution, because $B(s_{2})-B_{c}(s_{2})$ is purely imaginary. So, the non-pole parts of $C_{3}$ and $D_{3}$ satisfy
@@
Due to principal value, the expansion of $C_{3}^{NP}$ in $\epsilon$ is subtle. When we calculate $B(s_{0})/x_{p}$, we encounter following quantity
$I_{0}(x)=\frac{(-x-i\epsilon)^{-\epsilon/2}}{x_{p}}.$ (96)
To extract its real part and imaginary part, we write $I_{0}(x)$ as
$I_{0}(x)=$ $\frac{|x|^{-\epsilon/2}}{x_{p}}[(-1-i\epsilon)^{-\epsilon/2}\theta(x)+\theta(-x)]$
$=$ $\frac{|x|^{-\epsilon/2}}{x_{p}}[\theta(x)+\theta(-x)+i\frac{\pi\epsilon}{2}\theta(x)+O(\epsilon^{2})]$
$=$ $\frac{|x|^{-\epsilon/2}}{x_{p}}[1+i\frac{\pi\epsilon}{2}\theta(x)+O(\epsilon^{2})].$ (97)
For the imaginary part, the expansion $|x|^{-\epsilon/2}=1-\frac{\epsilon}{2}\ln|x|+\cdots$ is not allowed, because we will encounter $\theta(x)/x_{p}$, which is meaningless. Actually, $\theta(x)x^{-\epsilon}/x_{p}$ should be understood as $\theta(x)x^{-1-\epsilon}$, which can be expanded as follows
$\theta(x)x^{-1-\epsilon/2}=\theta(x)\Big[-\frac{2}{\epsilon}\delta(x)+\frac{1}{(x)_{+}}-\frac{\epsilon}{2}\Big(\frac{\ln x}{x}\Big)_{+}+O(\epsilon^{2})\Big].$ (98)
Now, the singularity at $x=0$ is extracted. Plus functions in the above is standard and defined in the region $x\in(0,1)$. On the other hand, the real part can be expanded directly,
$\frac{|x|^{-\epsilon/2}}{x_{p}}=\frac{1}{x_{p}}(1-\frac{\epsilon}{2}\ln|x|+O(\epsilon^{2})).$ (99)
Without $\theta(x)$, $1/x_{p}$ and $\ln|x|/x_{p}$ are well defined. The final result is
$I_{0}(x)=$ $\frac{1}{x_{p}}\Big(1-\frac{\epsilon}{2}\ln|x|\Big)+i\pi\theta(x)\Big(-\delta(x)+\frac{\epsilon}{2}\frac{1}{(x)_{+}}\Big)+O(\epsilon^{2}).$ (100)
For definiteness, we let $\theta(0)=1$. Compared with the real part, the imaginary part of $I_{0}(x)$ is not suppressed by $\epsilon$.
After this point is clear, $C_{3}^{NP}$ can be obtained, i.e.,
$\frac{p_{a}^{+}}{N_{c}C_{F}}C_{3}^{NP}=$ $\frac{x_{2}}{4\pi s_{2}\Gamma(1-\frac{\epsilon}{2})}\Big(\frac{4\pi\mu^{2}}{s_{2}}\Big)^{\epsilon/2}\Big\{C_{A}\Big[-\frac{\theta(-x)}{x_{2}-x}+\delta(x)\Big(-\frac{2}{\epsilon}-2-\ln x_{2}\Big)+\frac{\theta(x)}{x_{+}}\Big]$
$+(C_{A}-2C_{F})\Big[-\frac{1}{x_{p}}-\frac{x_{2}\theta(x-x_{2})}{x^{2}}\Big(\frac{2}{\epsilon}+2\Big)-\frac{x_{2}\theta(x_{2}-x)}{(x^{2})_{p}}\ln\frac{x_{2}-x}{x_{2}}\Big]\Big\}.$ (101)
$x_{2}$ or $s_{2}$ is always positive.
@@
$D_{3}^{NP}$ can be obtained from $C_{3}^{NP}$ from eq.(95). Both $t_{a}^{\mu\nu}$ and $t_{b}^{\mu\nu}$ appear in this part, while only $t_{a}^{\mu\nu}$ appears in SGP part and in tree level result. Moreover, the coefficients here are divergent. Different from twist-2 cases, the new tensor structure with divergent coefficients does not imply a breaking of factorization, as we will explain later.
@@ V.2 Corrections to derivative part of $W^{\mu\nu}$
$H_{6}^{\mu\nu}=H_{4}^{\mu\nu}=-i\frac{\partial\delta^{n-2}(q_{\perp})}{\partial q_{\perp}^{\rho}}\tilde{s}_{\perp}^{\rho}Tr[H^{\mu\nu}\otimes T^{a}\gamma^{-}\otimes\gamma^{+}].$ (104)
@@
which is just the correction to quark form factor, as pointed out in [15]. Note that $s_{2}=Q^{2}$ for virtual correction.
@@ V.3 Total virtual corrections
Now, the complete virtual correction to hadronic tensor is the sum of eqs.(91,102,105), that is,
@@
If there is no non-pole contribution, the virtual correction has the same structure as tree level hadronic tensor. The pole contribution, the last second line of above result, satisfies QED gauge invariance, as shown for tree level hadronic tensor. The correction to derivative part is the same as quark form factor, in agreement with the conclusion of [15]. In [15], such correction is inferred from Ward identity for longitudinal gluon $G^{+}$. Here we recover the result by direct calculation. The non-pole contribution, the last line, also satisfies such invariance: with $q_{\perp}=0$,
@@
Using eq.(8), we have
$t_{a}^{\mu\nu}L_{\mu\nu}\Big|_{q_{\perp}=0}=\frac{4s_{2}}{x_{2}}\cos\theta\tilde{s}_{\perp}\cdot l_{\perp,cs},\quad t_{b}^{\mu\nu}L_{\mu\nu}\Big|_{q_{\perp}=0}=-\frac{4s_{2}}{x_{b}}\cos\theta\tilde{s}_{\perp}\cdot l_{\perp,cs},\quad g_{\perp}^{\mu\nu}\frac{\partial L_{\mu\nu}}{\partial q_{\perp}^{\rho}}\Big|_{q_{\perp}=0}=0.$ (109)
Then, $I\langle L\rangle$ and $I\langle P_{7}\rangle$ can be obtained easily, which are written as
@@
$I\langle L\rangle\Big|^{NP}_{v}=$ $2\times l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int\frac{d\xi}{\xi}\frac{dx_{b}}{x_{b}}\delta(1-\hat{x}_{b})\bar{q}(x_{b})\int_{x^{*}_{2}-1}^{1}dx_{2}\Big[$
$N_{c}^{2}(\frac{128}{\epsilon}+128+64\ln x^{*}_{2})\delta(x_{2})T_{F}(x^{*}_{2}-x_{2},x^{*}_{2})$
$+64T_{F}(x^{*}_{2}-x_{2},x^{*}_{2})\Big((\frac{2}{\epsilon}+2)\frac{x^{*}_{2}\theta(x_{2}-x^{*}_{2})}{x_{2}^{2}}+\frac{N_{c}^{2}\theta(-x_{2})}{x^{*}_{2}-x_{2}}-\frac{N_{c}^{2}\theta(x_{2})}{(x_{2})_{+}}+\frac{1}{(x_{2})_{p}}+\frac{x^{*}_{2}\theta(x^{*}_{2}-x_{2})}{(x_{2}^{2})_{p}}\ln\frac{x^{*}_{2}-x_{2}}{x^{*}_{2}}\Big)\Big].$ (112)
@@
$\hat{x}_{2}\equiv\frac{x_{2}^{*}}{x_{2}},\ \hat{x}_{b}\equiv\frac{x_{b}^{*}}{x_{b}},\ x_{2}^{*}=\xi,\ x_{b}^{*}=\tau_{\xi},\ \xi=\frac{q^{+}}{p_{a}^{+}},\ \tau=\frac{Q^{2}}{2p_{a}\cdot p_{b}},\ \tau_{\xi}=\frac{\tau}{\xi}.$ (114)
@@
Recently, [21] gives NLO corrections to SSA in lepton-hadron collision. For virtual corrections, their diagrams are the same as ours. Their result (eq.(37) of [21]) contains both SGP contribution and non-pole contribution. Our SGP correction is the same as theirs, up to a $\pi^{2}$, which is due to the continuation of $Q^{2}$. Their result contains two kinds of non-pole contributions. In their notation, the first one is for the region $x^{\prime}>x$, which corresponds to the first line of our eq.(101). The second nonpole contribution is for the region $x^{\prime}<0$, which corresponds to the second line of our eq.(101). They found that the divergent parts of these two non-pole contributions can be subtracted out by the evolution kernel of twist-3 distribution function. This is also the case of our results. We will show this later. It should be noted that [21] performs the calculation in light-cone gauge, but we use Feynman gauge. So, they use $G_{\perp}$ to do collinear expansion, but we use $G^{+}$. Moreover, the integration order is different. [21] first integrates out the loop momentum and then do expansion in $k_{T}$, with $k_{T}$ the small transverse momentum of initial quark. But we do expansion in transverse momentum first, and then reduce the resulting integrals. Reproducing the structure of the virtual correction of [21] is a support to our calculation scheme.
@@ VI Real corrections
For real corrections, $q_{\perp}\neq 0$ is assumed. $W^{\mu\nu}$ thus depends on $p_{a}^{\mu},p_{b}^{\mu},q_{\perp}^{\mu}$ and $\tilde{s}_{\perp}^{\mu}$. Because of QED gauge invariance $q_{\mu}W^{\mu\nu}=0$, there are six independent tensor structures for transversely polarized case[35]. For our convenience, the tensors can be chosen as
@@
The dot product represents the contraction of Lorentz indices, e.g., $P_{j}\cdot W=P_{j,\mu\nu}W^{\mu\nu}$. Eq.(117) can be solved directly. But the coefficients before $P_{i}\cdot W$ may depend on spin vector $\tilde{s}_{\perp}$ in a very complicated way, since $P_{i}\cdot P_{j}$ depends on $\tilde{s}_{\perp}$. Noticing that $W_{i}$ does not depend on the direction of $q_{\perp}$, we can integrate out the angles of $q_{\perp}$ on both sides of eq.(117) and then solve the obtained equations. According to this treatment, we successfully get
$W_{i}=\frac{2-\epsilon}{\Omega_{n-2}\tilde{s}_{\perp}^{2}}\sum_{j}C_{ij}\xi_{j},\ \xi_{j}\equiv\int d\Omega_{n-2}P_{j}\cdot W,$ (118)
@@
Notice that $C_{ij}$ is symmetric, $C_{ij}=C_{ji}$ and $q_{t}=|\vec{q}_{\perp}|$.
@@
$=$ $\sum_{i,j}\int d^{n}q\delta(q^{2}-Q^{2})L\cdot P_{i}\frac{2-\epsilon}{\Omega_{n-2}\tilde{s}_{\perp}^{2}}C_{ij}\xi_{j}.$ (120)
@@
$\sum_{i,j}\Big(\int d\Omega_{n-2}L\cdot P_{i}\Big)\frac{2-\epsilon}{\Omega_{n-2}\tilde{s}_{\perp}^{2}}C_{ij}\xi_{j}=l_{\perp,cs}\cdot\tilde{s}_{\perp}\sum_{i}a_{i}\xi_{i}\frac{1}{\tilde{s}_{\perp}^{2}}.$ (123)
@@
This is our main formula for real corrections. There is no $\tilde{s}_{\perp}$ or $l_{\perp,cs}$ in $\tilde{W}_{i}$ now. With $W^{\mu\nu}$ replaced by $\tilde{W}_{i}$, the formula eq.(41) can be applied. Further calculation is the same as that for real correction of weighted cross section studied in [15].
Same as weighted cross section, there is no contribution from two-point correlation functions, i.e.,Fig.3. The amplitudes in these diagrams do not contain any absorptive part. So, $M^{(0)}_{\gamma_{\perp}\gamma_{5}}$ does not contribute and we just need to consider three-point distribution functions. Still, three types of poles contribute: SGP, SFP and hard pole(HP). The details can be found in [15], here we just present the final result for each pole contribution. We note that for HP and SFP, the momentum fraction of initial gluon $x\neq 0$, and it can be shown easily that the collinear expansion based on $G^{+}$ or $G_{\perp}$ leads to the same hard coefficients[7]. We have checked this by direct calculation. In App.B, we present some details for the calculation of HP contribution based on $G_{\perp}$ expansion. The procedure based on $G^{+}$ expansion is too lengthy and not shown. For SGP contribution, we use $G^{+}$ to do collinear expansion only. Especially, for Fig.4, the contribution from $M^{(1)}_{\gamma_{\perp}\gamma_{5}}$ vanishes, because corresponding $W^{\mu\nu}$ projected by $\gamma_{5}\gamma_{\perp}$ is anti-symmetric in $\mu,\nu$. $M^{(1)}_{\gamma_{\perp}}$ does not contribute because
@@
which can be shown from PT symmetry or from the second equation of eq.(26) and eq.(32). The SGP contribution from $M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k^{+},k_{1}^{+})$ also vanishes, although $M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k^{+},k_{1}^{+})$ may be nonzero at $k^{+}=0$. In this work, we consider only the contribution relevant to $\bar{q}\otimes T_{F}$. Because $M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}$ is related to $T_{\Delta}$ rather than $T_{F}$, we do not consider the contribution from this matrix elements in the following. Our SGP contributions are given by $M^{(1)}_{\gamma^{+},\partial_{\perp}\psi}$ and $M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}$. As illustrated in Sec.III, SGP contributions from these two matrix elements can be expressed by $T_{F}$. In addition, we also use $L^{\mu\nu}$ instead of the projection operators $P_{i}^{\mu\nu}$ to calculate. The same $I\langle L\rangle$ is obtained. This is a check of our calculation. In the calculation, we first integrate out $k_{g}$, then do expansion in $k_{\perp}$ and $k_{2\perp}$. After this step, $q_{\perp}$ is integrated out. It is easy to see if we first do expansion in $k_{\perp}$ and $k_{2\perp}$, the result is the same. Such an order of expansion and integration is different from the order taken in [21], where $k_{\perp}$ or $k_{2\perp}$ expansion is done after phase space integration is completed.
@@
VI.1 SGP contributions
SGP contribution is given by Fig.4. The short bar indicates the propagator is on-shell. Fig.4(c,d) are mirror diagrams[7], which do not contribute. Our calculation confirms this.